A sales representative of a company using a personal car receives $180 per day for lodging and meals plus $0.70 per mile driven. Write a linear equation giving the daily cost C to the company in terms of x, the number of miles driven.
1
Expert's answer
2019-09-13T10:03:56-0400

Let C(x) be our linear equation. C(x) = k*x + b. If we drive 0 miles, the cost is:

C(0) = k*0 + b = 180, so b = 180.

k is 0.7 - it's obvious from the task. So C(x) = 0.7*x + 180.

Answer: C(x) = 0.7*x + 180.